### v1

@@ Abstract
keywords:
Item response, Latent class model, Misclassification rate, Spectral clustering
@@ 1 Introduction
Subgroup analysis, the analysis of heterogeneous treatment effects across multiple sub-populations, plays a crucial role in many scientific fields. In biomedical studies, modern therapies are often found to be helpful in some subpopulations but much less so for others. For example, amado2008wild noted that panitumumab shows heterogeneous treatment effects in metastatic colorectal cancer patients with different subtypes of the KRAS gene. When this happens, subgroup analysis can enhance our understanding of the treatment and provide guidance on where the treatment should be implemented. Besides biomedical studies, subgroup analysis is also widely used in the broad areas of data science, such as educational studies (yin2024effects) and policy learning (shi2021online).
@@
How to make valid and efficient estimation and inference on the treatment effects of latent subgroups is a critical question in the broad areas of data science, such as educational, social, and health studies (page2015principal). One motivating example of this paper is Trends in International Mathematics and Science Study (TIMSS), a large-scale educational assessment that investigates the students’ performance in mathematics and science around the world (martin2015timss). In TIMSS, one educational question of interest is how private science lessons impact the performance of students with different motivations in science assessment (lyu2023estimating). Although student motivation is captured by about 70 binary item responses (agree or disagree) with respect to opinions on learning mathematics and sciences from TIMSS, defining $2^{70}$ classical subgroups directly based on answers to these questions is clearly neither infeasible nor interpretable. Indeed, students with different motivations are often viewed as latent subgroups inferred from imperfect measurements, the answers to questionnaire items regarding opinions on learning mathematics and sciences (bechter2018student). Moreover, TIMSS is an observational study with potentially high-dimensional confounders, such as family socioeconomic status and parent education. Such problem setups are not unique to TIMSS and also arise in other scientific disciplines, such as social science surveys and biomedical studies. For example, in studies of National Longitudinal Survey of Adolescent Health, adolescents with different risks are viewed as latent subgroups inferred from 6 binary risk characteristics (lanza2013alatent), and in studies of acute respiratory distress syndrome (ARDS) associated with COVID-19, ARDS subphenotypes are viewed as latent subgroups inferred from demographics, vital signs and laboratory values in electronic health records (sinha2021latent). Both survey and electronic health records are observational data with potentially high-dimensional confounders, such as school management and use of immunomodulatory drugs. Therefore, it is desirable to propose appropriate methods to accurately and efficiently estimate and infer the treatment effects across latent subgroups while adjusting for confounders.
@@
Latent subgroup analysis with observational data can be much more challenging than that for classical subgroups as it typically consists of two inter-connected models (zhou2022subgroup), a latent class model and a subgroup outcome model. The latent class model relates latent subgroup memberships with multivariate imperfect measurements, such as item responses collected from assessments or surveys. The subgroup outcome model evaluates the treatment effects in latent subgroups by modeling the relationship between latent subgroup memberships, treatment, potential confounders, and outcomes of interest. To estimate and infer latent subgroup effects, some existing methods consider the classical one-stage framework, which fits two models simultaneously by maximizing the (penalized) joint likelihood (shen2020subgroup; elliott2020methods; kang2023identification). The one-stage framework is widely adopted in practice, but it is likelihood-based and might encounter several issues with high-dimensional observational data. First, the one-stage framework could be computationally infeasible or inefficient when many potential confounders or item responses are included, leading to a complex joint likelihood function and challenging optimization problems (schuler2014addressing). Second, the one-stage framework might be theoretically biased when the joint likelihood is misspecified or the parameter space is non-identifiable or increasingly complex with high-dimensional potential confounders (vermunt2010latent; shen2015inference; lythgoe2019latent). Third, the one-stage framework fits the two models simultaneously, in conflict with the common belief of practitioners that subgroup effects should be evaluated after determining subgroup memberships to bear a clear interpretation of subgroups not subject to the outcomes or confounders. Such a one-stage fitting would also increase the risk of data snooping and become infeasible when the data are stored separately (vermunt2010latent).
@@
Besides the one-stage framework, there is also considerable research dedicated to the two-stage framework which fits two models separately by first estimating the latent class model and then evaluating subgroup effects with estimated latent memberships (bakk2021relating; kang2020causal). The two-stage framework is intuitive and easy to implement, might accommodate many potential confounders, and has been widely adopted in the analysis of latent subgroups with observational data. Examples include the study of latent subgroup effects of preschool programs to disadvantaged children (cooper2014benefits) and corticosteroids to Covid-related ARDS patients (sinha2021latent). However, as the estimated latent membership might not be the ground truth, the two-stage framework might introduce misclassification in the first stage and bias the analysis of subgroup effects in the second stage. To better support the practical use of the two-stage framework, it is desirable to understand when and how the two-stage latent subgroup analysis remains valid with observational data.
@@
Existing two-stage approaches for latent subgroups typically fit the latent class model via the EM algorithm (bakk2014relating), which is computationally inefficient and lacks theoretical guarantees for the latent class model with many items (kwon2019global). Although several approaches have been proposed to correct the misclassification bias in the two-stage framework (bolck2004estimating; bakk2018two), they are typically designed for point estimation in the absence of covariates and are not directly applicable for inference on the subgroup treatment effects from observational data. To the best of our knowledge, rigorous studies of misclassification bias and efficient and provable two-stage approaches for latent subgroups are still lacking especially in observational studies, and we aim to bridge the gap in this paper.
@@ 2.1 Problem Setting
To start with, we model the outcome and relate $Y$ with $X$, $D$, and $Z$. In particular, following the subgroup analysis setup in guo2023assessing and belloni2014inference, we assume $Y$ can be approximated by a linear transformation of $X$ and consider a linear subgroup outcome model
@@
where $\alpha_{g}$, $\mu_{g}$ and $\mathbf{\beta}$ are the regression coefficients and $\varepsilon$ is a Gaussian error term. Under the strong ignorability assumption $\{Y(1),Y(0)\}\bot D\mid(Z,X)$, we have $\mu_{g}=\mathbb{E}[Y(1)-Y(0)\mid Z=g]$ is the treatment effect of the $g$-th latent subgroup and $\mu=(\mu_{1},\ldots,\mu_{G})^{\mathsf{T}}$ is thus the quantity of interest. When the subgroup membership $Z$ is observed, estimation and inference of the subgroup effects $\mu$ have been discussed in imai2013estimating and guo2023assessing. However, here, we focus on a scenario where $Z$ is latent and unobserved.
Next, we model the latent subgroup membership and relate $Z$ with $R$. In particular, we consider a latent class model with binary item responses (goodman1974exploratory; clogg1995latent) where the latent class variable $Z$ follows the categorical distribution with proportion parameters $\tau=\{\tau_{g}\}_{g=1}^{G}$; i.e., $\mathbb{P}(Z=g)=\tau_{g}$, $g=1,\ldots,G$, and given $Z=g$, the item responses are assumed to be conditionally independent Bernoulli random variables with item parameters $\theta_{g}=(\theta_{1,g},\ldots,\theta_{J,g})^{\mathsf{T}}$, $g\in[G]$; i.e.,
@@
where $r_{j}$ is the $j$-th element of $r$. This modeling is different from that in fusion-type subgroup analysis (ma2017concave) where subgroup membership is not related with item responses. We denote the item response matrix by $\mathbf{R}=\{R_{i}\}_{i=1}^{n}\in\{0,1\}^{n\times J}$ and the item parameter matrix by $\boldsymbol{\Theta}=\{\theta_{j,g}\}^{J\times G}\in[0,1]^{J\times G}$.
@@
Note that in Eqs. (1) and (2), we do not model the relationship between the potential confounders $X$ and the latent subgroup membership $Z$ or item responses $R$. Such flexibility allows the distribution of potential confounders to be heterogeneous across different subgroups and more importantly, a variable can serve the roles of both a confounder and an item response. In other words, $X$ can include item responses $R$. In addition, we consider a linear subgroup outcome model in Eq. (1) and a latent class model in Eq. (2) for the sake of simplicity. Extensions to generalized linear models (beyond linear models) and other mixture models (beyond latent class model) are feasible, as long as: (a) the confounding effect can be well approximated by a (sparse) linear combination of potential confounders as discussed in wang2019debiased and guo2023assessing, and (b) the response data in the mixture model are sub-Gaussian as described in zhang2022leave.
@@
Figure 3(a)LABEL:fig:model illustrates the graphical model representation of latent subgroup analysis with observational data, which consists of two inter-connected models: the latent class model and the subgroup outcome model. To estimate and infer the latent subgroup effect $\mu$, one can consider either a one-stage or a two-stage framework. The one-stage framework, as illustrated by Figure 3(a)LABEL:fig:one_stage, maximizes the joint likelihood $\mathbb{P}(Y,D,R,X)$ of all variables and fits the two models simultaneously. Although the validity and practical implementation of the one-stage framework have been studied for classical low-dimensional settings, it typically relies on the correct specification of $\mathbb{P}(Y,D,R,X)$ and the low statistical complexity of the whole parameter space $(\tau,\boldsymbol{\Theta},\alpha,\mu,\beta)$. In observational studies, such requirements are hard to satisfy as the number of potential confounders plus item responses $J+p$ is often large and the joint likelihood of the two models $\mathbb{P}(Y,D,R,X)$ is often complex with numerous parameters. Moreover, simultaneously fitting two models contradicts the usual practice in subgroup analysis that subgroup effects $\mu$ are evaluated after subgroup memberships have been determined. This might even be impossible when the data for item responses $R$ and the others $(Y,D,X)$ are stored separately. To address the above issues, the two-stage framework, as illustrated by Figure 3(a)LABEL:fig:two_stage, has been considered which proceeds by first estimating latent subgroup membership $\widehat{Z}$ based on the item response $R$ alone and then evaluating the latent subgroup effect $\mu$ by regressing $Y$ on $\widehat{Z},D\widehat{Z},X$ (bakk2021relating).
@@
((a)) Illustration of two frameworks for latent subgroup analysis with observational data.
@@
Obviating the need for jointly modeling all variables and simultaneously fitting two models, the two-stage framework provides a potentially more feasible way to conduct latent subgroup analysis with observational data. However, we must note that the estimated memberships $\widehat{Z}$ are not necessarily equal to the true $Z$ and such misclassification might introduce bias. As illustrated by the numerical example in Figure 3, misclassification bias might not diminish by simply considering different fitting methods of the latent class model, such as hard assignment (nagin2005group) and soft assignment (goodman20071), or increasing the sample size (neyman1948consistent). This is because even when we know the true parameter of the latent class model, a subject might be misclassified with non-zero probability; i.e., $P(\widehat{Z}=g|Z=g)=\sum_{r\in\mathcal{R}}P(\widehat{Z}=g|R=r)P(R=r|Z=g)<1$, unless $P(\widehat{Z}=g|R=r^{*})=1$ for any $r^{*}\in\mathcal{R}$. Such intrinsic uncertainty of subgroup assignments in the first stage might substantially bias the analysis of subgroup effects in the second stage, and it is unclear whether and when the misclassification bias is negligible.
@@ 2.2 Maximum Misclassification Rate
Directly studying the maximum misclassification rate is challenging due to the self-correlation issue induced by $\widehat{Z}$. In particular, neither $\widehat{Z}_{i}$ nor $\widehat{Z}_{i}-Z_{i}$ can be treated as i.i.d. random variables, and indeed, their distributions depend on how the latent class model is fitted and the misclassification bias can be very complicated in the presence of potential confounders. Therefore, classical asymptotic theories and error-in-variable models are not directly applicable here (schennach2016recent). To derive the maximum misclassification rate, we adopt appropriate bridges between $\widehat{\mu}$ and $\mu$ and quantify the impacts of misclassified subjects sharply in both low and high dimensions.
@@
In the high-dimensional setting where $p+2G>n$ and $p$ and $G$ might increase with $n$, Lasso and debiased Lasso, $\widehat{\mu}_{\text{lasso}}$ in Eq. (5) and $\widehat{\mu}_{\text{dl}}$ in Eq. (6) and $\widehat{\mathbf{\Sigma}}_{\text{dl}}$, are commonly used in fitting the subgroup outcome model for estimation (Lasso) and inference (debiased Lasso) on the latent subgroup effects $\mu$. In high dimensions, we can not adopt the oracle regularized estimates based on $Z$, $\widetilde{\mu}_{\text{lasso}}$ and $\widetilde{\mu}_{\text{dl}}$, as bridges directly, as they often lack explicit forms (hastie2015statistical). Instead, we adopt the oracle regularized loss function based on $Z$ as a bridge and bound the difference between the regularized loss function based on $Z$ and that based on $\widehat{Z}$ by the misclassification rate as detailed in Lemma 5 in Appendix A. We assume classical regularity assumptions, Assumptions A.2-A.4 in Appendix A, that guarantee the validity of $\widetilde{\mu}_{\text{lasso}}$ and $\widetilde{\mu}_{\text{dl}}$ in high dimensions. Then, Theorem 2 states that the two-stage Lasso $\widehat{\mu}_{\text{lasso}}$ is consistent with oracle rate and the two-stage debiased Lasso $\widehat{\mu}_{\text{dl}}$ leads to valid confidence intervals with $\widehat{\mathbf{\Sigma}}_{\text{dl}}$ if the misclassification rate converges to 0 at a rate not slower than $\sqrt{\log(p+2G)/n}$ and faster than $1/n$ respectively. Both rates are tighter than their counterparts in low dimensions to compensate for the more challenging task in high dimensions. The rate requirement for estimation is more relaxed when there are more subgroups or confounders as in that case, the oracle rate is also slower. Similar to the low-dimensional setting, the rate requirement for inference is tighter than that for estimation, and any misclassification rate asymptotically larger than the desired rates will lead to non-oracle estimation or invalid inference on $\mu$ in any compact parameter space subject to the construction of compatibility condition (Assumption A.2 in Appendix A), a classical and almost necessary condition required in high dimensions as discussed in van2017some. It is in this sense that we call the desired rate the maximum misclassification rate in high dimensions.

### v2

@@ Abstract
Keywords: Item response; Latent class model; Misclassification rate; Spectral clustering.
@@ 1 Introduction
Subgroup analysis, the analysis of heterogeneous treatment effects across multiple sub-populations, plays a crucial role in many scientific fields. In biomedical studies, modern therapies are often found to be helpful in some subpopulations but much less so for others. For example, Amado et al., (2008) noted that panitumumab shows heterogeneous treatment effects in metastatic colorectal cancer patients with different subtypes of the KRAS gene. When this happens, subgroup analysis can enhance our understanding of the treatment and provide guidance on where the treatment should be implemented. Besides biomedical studies, subgroup analysis is also widely used in the broad areas of data science, such as educational studies (Yin et al.,, 2024) and policy learning (Shi et al.,, 2021).
@@
How to make valid and efficient estimation and inference on the treatment effects of latent subgroups is a critical question in the broad areas of data science, such as educational, social, and health studies (Page et al.,, 2015). One motivating example of this paper is Trends in International Mathematics and Science Study (TIMSS), a large-scale educational assessment that investigates the students’ performance in mathematics and science around the world (Martin et al.,, 2015). In TIMSS, one educational question of interest is how private science lessons impact the performance of students with different motivations in science assessment (Lyu et al.,, 2023). Although student motivation is captured by about 70 binary item responses (agree or disagree) with respect to opinions on learning mathematics and sciences from TIMSS, defining $2^{70}$ classical subgroups directly based on answers to these questions is clearly neither infeasible nor interpretable. Indeed, students with different motivations are often viewed as latent subgroups inferred from imperfect measurements, the answers to questionnaire items regarding opinions on learning mathematics and sciences (Bechter et al.,, 2018). Moreover, TIMSS is an observational study with potentially high-dimensional confounders, such as family socioeconomic status and parent education. Such problem setups are not unique to TIMSS and also arise in other scientific disciplines, such as social science surveys and biomedical studies. For example, in studies of National Longitudinal Survey of Adolescent Health, adolescents with different risks are viewed as latent subgroups inferred from 6 binary risk characteristics (Lanza and Rhoades,, 2013), and in studies of acute respiratory distress syndrome (ARDS) associated with COVID-19, ARDS subphenotypes are viewed as latent subgroups inferred from demographics, vital signs and laboratory values in electronic health records (Sinha et al.,, 2021). Both survey and electronic health records are observational data with potentially high-dimensional confounders, such as school management and use of immunomodulatory drugs. Therefore, it is desirable to propose appropriate methods to accurately and efficiently estimate and infer the treatment effects across latent subgroups while adjusting for confounders.
@@
Latent subgroup analysis with observational data can be much more challenging than that for classical subgroups as it typically consists of two inter-connected models (Zhou et al.,, 2022), a latent class model and a subgroup outcome model. The latent class model relates latent subgroup memberships with multivariate imperfect measurements, such as item responses collected from assessments or surveys. The subgroup outcome model evaluates the treatment effects in latent subgroups by modeling the relationship between latent subgroup memberships, treatment, potential confounders, and outcomes of interest. To estimate and infer latent subgroup effects, some existing methods consider the classical one-stage framework, which fits two models simultaneously by maximizing the (penalized) joint likelihood (Shen and Qu,, 2020; Elliott et al.,, 2020; Kang and Huang,, 2023). The one-stage framework is widely adopted in practice, but it is likelihood-based and might encounter several issues with high-dimensional observational data. First, the one-stage framework could be computationally infeasible or inefficient when many potential confounders or item responses are included, leading to a complex joint likelihood function and challenging optimization problems (Schuler et al.,, 2014). Second, the one-stage framework might be theoretically biased when the joint likelihood is misspecified or the parameter space is non-identifiable or increasingly complex with high-dimensional potential confounders (Vermunt,, 2010; Shen and He,, 2015; Lythgoe et al.,, 2019). Third, the one-stage framework fits the two models simultaneously, in conflict with the common belief of practitioners that subgroup effects should be evaluated after determining subgroup memberships to bear a clear interpretation of subgroups not subject to the outcomes or confounders. Such a one-stage fitting would also increase the risk of data snooping and become infeasible when the data are stored separately (Vermunt,, 2010).
@@
Besides the one-stage framework, there is also considerable research dedicated to the two-stage framework which fits two models separately by first estimating the latent class model and then evaluating subgroup effects with estimated latent memberships (Bakk and Kuha,, 2021; Kang et al.,, 2020). The two-stage framework is intuitive and easy to implement, might accommodate many potential confounders, and has been widely adopted in the analysis of latent subgroups with observational data. Examples include the study of latent subgroup effects of preschool programs to disadvantaged children (Cooper and Lanza,, 2014) and corticosteroids to Covid-related ARDS patients (Sinha et al.,, 2021). However, as the estimated latent membership might not be the ground truth, the two-stage framework might introduce misclassification in the first stage and bias the analysis of subgroup effects in the second stage. To better support the practical use of the two-stage framework, it is desirable to understand when and how the two-stage latent subgroup analysis remains valid with observational data.
@@
Existing two-stage approaches for latent subgroups typically fit the latent class model via the EM algorithm (Bakk et al.,, 2014), which is computationally inefficient and lacks theoretical guarantees for the latent class model with many items (Kwon et al.,, 2019). Although several approaches have been proposed to correct the misclassification bias in the two-stage framework (Bolck et al.,, 2004; Bakk and Kuha,, 2018), they are typically designed for point estimation in the absence of covariates and are not directly applicable for inference on the subgroup treatment effects from observational data. To the best of our knowledge, rigorous studies of misclassification bias and efficient and provable two-stage approaches for latent subgroups are still lacking especially in observational studies, and we aim to bridge the gap in this paper.
@@ 2.1 Problem Setting
To start with, we model the outcome and relate $Y$ with $X$, $D$, and $Z$. In particular, following the subgroup analysis setup in Guo et al., (2023) and Belloni et al., (2014), we assume $Y$ can be approximated by a linear transformation of $X$ and consider a linear subgroup outcome model
@@
where $\alpha_{g}$, $\mu_{g}$ and $\mathbf{\beta}$ are the regression coefficients and $\varepsilon$ is a Gaussian error term. Under the strong ignorability assumption $\{Y(1),Y(0)\}\bot D\mid(Z,X)$, we have $\mu_{g}=\mathbb{E}[Y(1)-Y(0)\mid Z=g]$ is the treatment effect of the $g$-th latent subgroup and $\mu=(\mu_{1},\ldots,\mu_{G})^{\mathsf{T}}$ is thus the quantity of interest. When the subgroup membership $Z$ is observed, estimation and inference of the subgroup effects $\mu$ have been discussed in Imai and Ratkovic, (2013) and Guo et al., (2023). However, here, we focus on a scenario where $Z$ is latent and unobserved.
Next, we model the latent subgroup membership and relate $Z$ with $R$. In particular, we consider a latent class model with binary item responses (Goodman,, 1974; Clogg,, 1995) where the latent class variable $Z$ follows the categorical distribution with proportion parameters $\tau=\{\tau_{g}\}_{g=1}^{G}$; i.e., $\mathbb{P}(Z=g)=\tau_{g}$, $g=1,\ldots,G$, and given $Z=g$, the item responses are assumed to be conditionally independent Bernoulli random variables with item parameters $\theta_{g}=(\theta_{1,g},\ldots,\theta_{J,g})^{\mathsf{T}}$, $g\in[G]$; i.e.,
@@
where $r_{j}$ is the $j$-th element of $r$. This modeling is different from that in fusion-type subgroup analysis (Ma and Huang,, 2017) where subgroup membership is not related with item responses. We denote the item response matrix by $\mathbf{R}=\{R_{i}\}_{i=1}^{n}\in\{0,1\}^{n\times J}$ and the item parameter matrix by $\boldsymbol{\Theta}=\{\theta_{j,g}\}^{J\times G}\in[0,1]^{J\times G}$.
@@
Note that in Eqs. (1) and (2), we do not model the relationship between the potential confounders $X$ and the latent subgroup membership $Z$ or item responses $R$. Such flexibility allows the distribution of potential confounders to be heterogeneous across different subgroups and more importantly, a variable can serve the roles of both a confounder and an item response. In other words, $X$ can include item responses $R$. In addition, we consider a linear subgroup outcome model in Eq. (1) and a latent class model in Eq. (2) for the sake of simplicity. Extensions to generalized linear models (beyond linear models) and other mixture models (beyond latent class model) are feasible, as long as: (a) the confounding effect can be well approximated by a (sparse) linear combination of potential confounders as discussed in Wang et al., (2019) and Guo et al., (2023), and (b) the response data in the mixture model are sub-Gaussian as described in Zhang and Zhou, (2024).
@@
Figure undef??LABEL:fig:model illustrates the graphical model representation of latent subgroup analysis with observational data, which consists of two inter-connected models: the latent class model and the subgroup outcome model. To estimate and infer the latent subgroup effect $\mu$, one can consider either a one-stage or a two-stage framework. The one-stage framework, as illustrated by Figure undef??LABEL:fig:one_stage, maximizes the joint likelihood $\mathbb{P}(Y,D,R,X)$ of all variables and fits the two models simultaneously. Although the validity and practical implementation of the one-stage framework have been studied for classical low-dimensional settings, it typically relies on the correct specification of $\mathbb{P}(Y,D,R,X)$ and the low statistical complexity of the whole parameter space $(\tau,\boldsymbol{\Theta},\alpha,\mu,\beta)$. In observational studies, such requirements are hard to satisfy as the number of potential confounders plus item responses $J+p$ is often large and the joint likelihood of the two models $\mathbb{P}(Y,D,R,X)$ is often complex with numerous parameters. Moreover, simultaneously fitting two models contradicts the usual practice in subgroup analysis that subgroup effects $\mu$ are evaluated after subgroup memberships have been determined. This might even be impossible when the data for item responses $R$ and the others $(Y,D,X)$ are stored separately. To address the above issues, the two-stage framework, as illustrated by Figure undef??LABEL:fig:two_stage, has been considered which proceeds by first estimating latent subgroup membership $\widehat{Z}$ based on the item response $R$ alone and then evaluating the latent subgroup effect $\mu$ by regressing $Y$ on $\widehat{Z},D\widehat{Z},X$ (Bakk and Kuha,, 2021).
@@
(a) Illustration of two frameworks for latent subgroup analysis with observational data.
@@
Obviating the need for jointly modeling all variables and simultaneously fitting two models, the two-stage framework provides a potentially more feasible way to conduct latent subgroup analysis with observational data. However, we must note that the estimated memberships $\widehat{Z}$ are not necessarily equal to the true $Z$ and such misclassification might introduce bias. As illustrated by the numerical example in Figure 3, misclassification bias might not diminish by simply considering different fitting methods of the latent class model, such as hard assignment (Nagin,, 2005) and soft assignment (Goodman,, 2007), or increasing the sample size (Neyman and Scott,, 1948). This is because even when we know the true parameter of the latent class model, a subject might be misclassified with non-zero probability; i.e., $P(\widehat{Z}=g|Z=g)=\sum_{r\in\mathcal{R}}P(\widehat{Z}=g|R=r)P(R=r|Z=g)<1$, unless $P(\widehat{Z}=g|R=r^{*})=1$ for any $r^{*}\in\mathcal{R}$. Such intrinsic uncertainty of subgroup assignments in the first stage might substantially bias the analysis of subgroup effects in the second stage, and it is unclear whether and when the misclassification bias is negligible.
@@ 2.2 Maximum Misclassification Rate
Directly studying the maximum misclassification rate is challenging due to the self-correlation issue induced by $\widehat{Z}$. In particular, neither $\widehat{Z}_{i}$ nor $\widehat{Z}_{i}-Z_{i}$ can be treated as i.i.d. random variables, and indeed, their distributions depend on how the latent class model is fitted and the misclassification bias can be very complicated in the presence of potential confounders. Therefore, classical asymptotic theories and error-in-variable models are not directly applicable here (Schennach,, 2016). To derive the maximum misclassification rate, we adopt appropriate bridges between $\widehat{\mu}$ and $\mu$ and quantify the impacts of misclassified subjects sharply in both low and high dimensions.
@@
In the high-dimensional setting where $p+2G>n$ and $p$ and $G$ might increase with $n$, Lasso and debiased Lasso, $\widehat{\mu}_{\text{lasso}}$ in Eq. (5) and $\widehat{\mu}_{\text{dl}}$ in Eq. (6) and $\widehat{\mathbf{\Sigma}}_{\text{dl}}$, are commonly used in fitting the subgroup outcome model for estimation (Lasso) and inference (debiased Lasso) on the latent subgroup effects $\mu$. In high dimensions, we can not adopt the oracle regularized estimates based on $Z$, $\widetilde{\mu}_{\text{lasso}}$ and $\widetilde{\mu}_{\text{dl}}$, as bridges directly, as they often lack explicit forms (Hastie et al.,, 2015). Instead, we adopt the oracle regularized loss function based on $Z$ as a bridge and bound the difference between the regularized loss function based on $Z$ and that based on $\widehat{Z}$ by the misclassification rate as detailed in Lemma 5 in Appendix A. We assume classical regularity assumptions, Assumptions A.2-A.4 in Appendix A, that guarantee the validity of $\widetilde{\mu}_{\text{lasso}}$ and $\widetilde{\mu}_{\text{dl}}$ in high dimensions. Then, Theorem 2 states that the two-stage Lasso $\widehat{\mu}_{\text{lasso}}$ is consistent with oracle rate and the two-stage debiased Lasso $\widehat{\mu}_{\text{dl}}$ leads to valid confidence intervals with $\widehat{\mathbf{\Sigma}}_{\text{dl}}$ if the misclassification rate converges to 0 at a rate not slower than $\sqrt{\log(p+2G)/n}$ and faster than $1/n$ respectively. Both rates are tighter than their counterparts in low dimensions to compensate for the more challenging task in high dimensions. The rate requirement for estimation is more relaxed when there are more subgroups or confounders as in that case, the oracle rate is also slower. Similar to the low-dimensional setting, the rate requirement for inference is tighter than that for estimation, and any misclassification rate asymptotically larger than the desired rates will lead to non-oracle estimation or invalid inference on $\mu$ in any compact parameter space subject to the construction of compatibility condition (Assumption A.2 in Appendix A), a classical and almost necessary condition required in high dimensions as discussed in van de Geer, (2017). It is in this sense that we call the desired rate the maximum misclassification rate in high dimensions.